OpenStudy (anonymous):

Five more then twice a number is -13. find the number

OpenStudy (anonymous):

Let number be x. 2x + 5 = -13 2x = -18 x = -9

OpenStudy (lilg132):

2x + 5 = -13 2x = -13 -5 2x = -18 2x / 2 = -18 / 2 x = -9

OpenStudy (anonymous):

thankyou!